A stock has an average return of 7% and a standard deviation of 8%. If returns are normally-distributed, what is the probability of an actual return: (a) above 15%; (b) below -9%; and (c) above 23%? no excel or charts pls handwork only
Answer #1

solution: Stock average retum = tolo Standard deviatim = 87. If retoms are normally distributed then probability of an actual= P(RC=-940) sp. (2<= (-91– 7.1.) 8-1.) = P(22=-2) =(2.2801] ©) above 2314 probability for the return above 234. = ENORM.pfst